# DeCo-SGD: Joint Optimization of Delay Staleness and Gradient Compression Ratio for Distributed SGD

### v1

@@ 3 Theoretical Analysis of DD-EF-SGD
We derive the convergence rate of DD-EF-SGD. Notation list and detailed proof are in the Appendix.
@@ 3.1 Regular Assumptions
We assume functions are $L$-smooth with the gradient noise of D-SGD presumed to exhibit zero mean and a variance of $\sigma^{2}$. We gauge data heterogeneity using constants $\zeta_{i}^{2}>0$, which cap the variance among the $n$ workers. In Theorem 2, we posit that objective functions are $\mu$-strongly convex. Detailed assumptions are shown below.
@@
Assumption 2 ($\mu$-strongly convexity). We assume $\mu$-strong convexity of $f_{i},i\in[n]$, that is for all $\mathbf{x},\textbf{y}\in\mathbb{R}^{d}$:
@@
$\mathbf{g}^{i}_{t}=\nabla f_{i}(\mathbf{x}_{t})+\boldsymbol{\xi}^{i},\quad\quad\quad\mathbb{E}_{\boldsymbol{\xi}^{i}}\boldsymbol{\xi}^{i}=\textbf{0}_{d},\quad\quad\quad\mathbb{E}_{\boldsymbol{\xi}^{i}}\lVert\boldsymbol{\xi}^{i}\rVert^{2}\leq\sigma^{2}.$ (3)
$\lVert\nabla f_{i}(\mathbf{x})\rVert^{2}\leq\zeta_{i}^{2}+\lVert\nabla f(\mathbf{x})\rVert^{2},\quad\zeta^{2}=\frac{1}{n}\sum_{i=1}^{n}\zeta_{i}^{2},\quad\forall\mathbf{x}\in\mathbb{R}^{d},i\in[n].$ (4)
These assumptions are standard, following previous works cui2022infocom ; stich2020communication . For Assumption 3, we can assume that the $\sigma^{2}$ increases with the gradient norm hardthreshold . In this way, by replacing $\sigma$ with a new $\sigma^{\prime}$, one can obtain a tighter bound. Similarly in Assumption 4, the general version is $\lVert\nabla f_{i}(\mathbf{x})\rVert^{2}\leq\zeta_{i}^{2}+Z\lVert\nabla f(\mathbf{x})\rVert^{2}$, and we take $Z=1$ by default. One can take a more general bound by adding $Z$. However, extending the analysis to cover these generalized assumptions is beyond the scope of this work.
3.2 Nested Virtual Sequences (NVS)
To solve the problem that the traditional Virtual Sequence method cannot analyze complex D-SGD variants like DD-EF-SGD. We propose NVS, which is composed of two parts. In the first part, we have
@@ 3.2 Nested Virtual Sequences (NVS)
Compared with Virtual Sequence, we change the second term in Eq. 5 from the update $\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$ into $\tilde{\mathbf{v}}_{t}$. We define the first virtual sequence $\tilde{\mathbf{x}}_{t}=\mathbf{x}_{t}-B_{t}$ and observe that $\tilde{\mathbf{x}}_{t+1}=\mathbf{x}_{t+1}-B_{t+1}=\tilde{\mathbf{x}}_{t}-\mathbf{v}_{t}-B_{t}-\tilde{\mathbf{v}}_{t}+\mathbf{v}_{t}=\tilde{\mathbf{x}}_{t}-\tilde{\mathbf{v}}_{t}$. In the second part, we use the virtual sequence framework based on the new items $\tilde{\mathbf{x}}_{t}$ and $\tilde{\mathbf{v}}_{t}$. We define the second noise term as $\tilde{B}_{t}$ and have:
$\tilde{\mathbf{x}}_{t+1}=\tilde{\mathbf{x}}_{t}-\tilde{\mathbf{v}}_{t},\quad\quad\quad\tilde{B}_{t+1}=\tilde{B}_{t}+\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}-\tilde{\mathbf{v}}_{t},\quad\quad\quad\hat{\mathbf{x}}_{t}=\tilde{\mathbf{x}}_{t}-\tilde{B}_{t}.$
@@
$\mathbf{v}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{C}_{\delta}(\mathbf{g}_{t-\tau}^{i}+e_{t-\tau}^{i}),\quad B_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}e_{t-\tau}^{i},$
$\tilde{\mathbf{v}}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i},\quad\tilde{B}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\sum_{j=1}^{\tau}\mathbf{g}_{t-j}^{i}.$ (6)
@@ 3.3 Convergence Rate of DD-EF-SGD
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth. There exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{4L\sqrt{\tau}},\frac{1}{4L\sqrt{\phi/\delta}}\}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\epsilon^{3/2}}+\frac{1}{\epsilon}+\frac{\sqrt{\tau}}{\epsilon}+\frac{\sqrt{\phi}}{\sqrt{\delta}\epsilon})\cdot L(f(\mathbf{x}_{0})-f^{*})\end{split}$ (7)
@@
($\phi$ determining the convergence in the non-degradation condition) In typical DML scenarios, $\frac{\tau}{n}\leq 1$. As the training scale increases, $\tau$ tends to decrease due to the larger model requiring larger computation time, and $n$ tends to increase, resulting in an even smaller $\frac{\tau}{n}$. While to reduce the communication traffic, $\delta$ tends to be $10\%$ or less, common in DML CocktailSGD ; grace , leading to the order of magnitude of $\phi$ larger than that of $\frac{\tau}{n}$. Then the convergence rate is $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}+\frac{\sqrt{\frac{\phi}{\delta}}+\sqrt{\tau}}{\epsilon})$. Similar to prior works dagctmc , the frist term $\frac{\sigma^{2}}{n\epsilon^{2}}$ is independent of $\delta$ and $\tau$. The second term $\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}$ has a greater impact on the overall convergence rate compared to the third term.
Let us focus on $\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}$, in LLMs pre-training scenarios, local datasets are typically allocated centrally, resulting in low data heterogeneity $\zeta$. Moreover, due to the small batch sizes commonly used in LLM training (limited by the constraint of GPU memory), $\sigma$ tends to be large. Under such circumstances, $\frac{\zeta^{2}}{\delta}\ll\sigma^{2}$ and $\phi$ captures the convergence rate. While in Federated Learning (large $\zeta$) or training small models (small $\sigma^{2}$), the key factor is $\phi^{\prime}=\frac{1-\delta}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}$. In this work, we focus on the first scenario where $\phi$ is the key factor that determines the convergence of DD-EF-SGD. For the analysis in Federated Learning or training small models, simply replacing $\phi$ with $\phi^{\prime}$ is enough and we do not discuss this later in the paper.
Remark 2.
(Degradation condition) DD-EF-SGD has two degradation conditions, namely, without compression ($\delta=1$) and without delayed aggregation ($\tau=0$). When $\delta=1$, $\phi=0$ and DD-EF-SGD degrades into DD-SGD. Then the convergence rate is equal to $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\frac{\tau}{n}\sigma^{2}}}{\epsilon^{3/2}}+\frac{\sqrt{\tau}}{\epsilon})$, same with the convergence rate of DD-SGD stich2022sharper ; stich2019error . When $\tau=0$, DD-EF-SGD degrades to D-EF-SGD, and $\phi=\frac{1-\delta}{\delta}$. Then the convergence rate becomes $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\zeta/\sqrt{\delta}+\sigma}{\sqrt{\delta}\epsilon^{3/2}}+\frac{L}{\delta\epsilon})$, same with the convergence rate of D-EF-SGD stich2020communication ; dagc .
Theorem 2 (Convex convergence rate of DD-EF-SGD).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-convex. Then there exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{8L\sqrt{\tau}},\frac{1}{16L\sqrt{\phi}}\}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L}{\mu}+\frac{\sqrt{L\tau}}{\mu}+\frac{\sqrt{L\phi}}{\mu})\end{split}$ (8)
iterations of DD-EF-SGD, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{T-1}\right\}$, selected probabilistically based on $(1-\min\left\{\frac{\mu\gamma}{2},\frac{\delta}{4}\right\})^{-t}$.
Remark 3.
Same to remarks in Theorem 1, in the degradation conditions, the convex convergence rate is the same as that of DD-SGD (under $\delta=1$) and D-EF-SGD (under $\tau=0$). In the non-degradation conditions, the convergence rate can be written into $\mathcal{O}(\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{\sqrt{\frac{L\phi}{\delta}}+\sqrt{L\tau}}{\mu})$, $\phi$ dominates the convergence rate of DD-EF-SGD in the convex case, too.

### v2

@@ 3 Theoretical Analysis of DD-EF-SGD
In this section, we establish the convergence guarantees for DD-EF-SGD. The key notations are summarized in Table 1. The training procedure of DD-EF-SGD is outlined in Algo. 1 Detailed proofs are provided in the Appendix (Sec. B.2B.3).
Table 1: Notation list.
Notation Description
$n$ number of workers
$f_{i}(\cdot)$ the local loss function of worker $i$
$\mathbf{x}^{i}_{t}$
the local model parameter in the $t$-th iteration
of the $i$-th worker
$\mathbf{x}_{t}$ the global model parameter in the $t$-th iteration
$\mathbf{g}^{i}_{t}$ the stochastic gradient of $\textbf{x}_{t}$ of the worker $i$
$\gamma$ the stepsize
$\mathbf{T}$ total number of iterations
$\delta$ compression ratio ($0<\delta\leq 1$)
$\mathbf{C}_{\delta}(\cdot)$
the sparsification compressor with
compression ratio $\delta$
$\tau$ the delayed staleness
$\mathbf{e}_{t}^{i}$
the local error term of the $i$-th worker
in the $t$-th iteration
$V_{g}$ the volume of the gradient (bits)
$E$ the frequency to update $\tau$ and $\delta$ with DeCo
$\bar{a}$
the average bandwidth (bits/s) of the slowest link
in previous $E$ iterations
$b$ the network end-to-end latency (s) of the slowest link
$T_{\text{comp}}$ computation time per iteration (s)
$T_{\text{avg}}$ average end-to-end training time per iteration (s)
$TS_{t}$ end time of the $t$-th computation (s)
$TC_{t}$ end time of the $t$-th communication (s)
Input: $n$, $\gamma$, $\mathbf{T}$, $C(\cdot)$, $\delta$, $\tau$
Output: $\textbf{x}_{\mathbf{T}}$
Initialize $\mathbf{x}_{0}$, $\mathbf{e}_{0}^{i}=\mathbf{0}_{d}$;
for $t\in[\mathbf{T}]$ do
 /* Worker side */
 for $i\in[n]$ do
      Worker $i$ waits until it receives $\textbf{x}_{t}$;
      Calculate $\mathbf{g}_{t}^{i}$ based on the global model $\mathbf{x}_{t}$;
    $\Delta_{t}^{i}=\textbf{C}_{\delta}(\textbf{e}_{t}^{i}+\gamma\textbf{g}_{t}^{i})$;
    $\textbf{e}_{t+1}^{i}=\textbf{e}_{t}^{i}+\gamma\textbf{g}_{t}^{i}-\Delta_{t}^{i}$;
      Send $\Delta_{t}^{i}$ to the server (non-blocking);
   end for
 /* Server side */
   Server waits until it receives $\Delta_{t-\tau}^{i}$ for all $i\in[n]$;
 $\textbf{x}_{t+1}=\textbf{x}_{t}-\frac{1}{n}\sum_{i=1}^{n}\Delta_{t-\tau}^{i}$;
   Broadcast $\textbf{x}_{t+1}$;
end for
Return $\textbf{x}_{\mathbf{T}}$;
Algorithm 1 $D$-EF-SGD
@@ 3.1 Regular Assumptions
We adopt standard assumptions, following previous works cui2022infocom; stich2020communication.
@@
$\mathbf{g}^{i}_{t}=\nabla f_{i}(\mathbf{x}_{t})+\boldsymbol{\xi}^{i},\quad\mathbb{E}_{\boldsymbol{\xi}^{i}}\boldsymbol{\xi}^{i}=\textbf{0}_{d},\quad\mathbb{E}_{\boldsymbol{\xi}^{i}}\lVert\boldsymbol{\xi}^{i}\rVert^{2}\leq\sigma^{2}.$ (2)
$\frac{1}{n}\sum_{i\in[n]}\lVert\nabla f_{i}(\mathbf{x})\rVert^{2}\leq\zeta^{2}+Z^{2}\lVert\nabla f(\mathbf{x})\rVert^{2},\forall\mathbf{x}\in\mathbb{R}^{d}.$ (3)
Under strongly convex objective functions, we use Assumption 4 in addition.
Assumption 4 ($\mu$-strongly convex). We assume $\mu$-strong convexity of $f_{i},i\in[n]$, that is for all $\mathbf{x},\textbf{y}\in\mathbb{R}^{d}$:
@@
3.2 Theoretical Tool: Nested Virtual Sequences (NVS)
A major challenge in analyzing DD-EF-SGD is the coupling effect between compression error and delay staleness. Traditional Virtual Sequence methods fail here because the noise term becomes recursively dependent on historical states, making it mathematically intractable.
We propose NVS, which is composed of two parts. In the first part, we have
@@ 3.2 Theoretical Tool: Nested Virtual Sequences (NVS)
Compared with Virtual Sequence, we change the second term in Eq. 5 from the update $\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$ into $\tilde{\mathbf{v}}_{t}$. We define the first virtual sequence $\tilde{\mathbf{x}}_{t}=\mathbf{x}_{t}-B_{t}$ and observe that $\tilde{\mathbf{x}}_{t+1}=\mathbf{x}_{t+1}-B_{t+1}=\tilde{\mathbf{x}}_{t}-\mathbf{v}_{t}-B_{t}-\tilde{\mathbf{v}}_{t}+\mathbf{v}_{t}=\tilde{\mathbf{x}}_{t}-\tilde{\mathbf{v}}_{t}$. In the second part, we use the virtual sequence framework based on the new items $\tilde{\mathbf{x}}_{t}$ and $\tilde{\mathbf{v}}_{t}$. We define the second noise term as $\tilde{B}_{t}$, having $\tilde{\mathbf{x}}_{t+1}=\tilde{\mathbf{x}}_{t}-\tilde{\mathbf{v}}_{t}$, $\tilde{B}_{t+1}=\tilde{B}_{t}+\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}-\tilde{\mathbf{v}}_{t}$ and $\hat{\mathbf{x}}_{t}=\tilde{\mathbf{x}}_{t}-\tilde{B}_{t}$.
@@
$\mathbf{v}_{t}$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbf{C}_{\delta}(\gamma\mathbf{g}_{t-\tau}^{i}+\mathbf{e}_{t-\tau}^{i}),$ $B_{t}$ $=\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{e}_{t-\tau}^{i},$ (6)
$\tilde{\mathbf{v}}_{t}$ $=\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i},$ $\tilde{B}_{t}$ $=\frac{\gamma}{n}\sum_{i=1}^{n}\sum_{j=1}^{\tau}\mathbf{g}_{t-j}^{i}.$ (7)
For form consistency in all iterative expressions, we define $\mathbf{g}_{t-\tau}^{i}=\mathbf{e}_{t-\tau}^{i}=\mathbf{0}^{d}$, for all $i\in[n]$, if $t-\tau\leq 0$. This means that when $t\leq\tau$, we have
$\mathbf{v}_{t}=\tilde{\mathbf{v}}_{t}=B_{t}=\mathbf{0}^{d},\quad\quad\tilde{B}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\sum_{j=1}^{t}\mathbf{g}_{t-j}^{i}.$
It is worth noting that since $\tau$ is typically much smaller than the total number of training iterations $T$, the case $t\leq\tau$ only occurs at the early stage of training. Consequently, it has a negligible influence on the overall convergence analysis and we will not discuss this boundary case.
@@ 3.3 Convergence Rate of DD-EF-SGD
Table 2: Summary of the convergence results for different algorithms under non-convex and $\mu$-strongly convex cases. We define $\phi=\frac{1-\delta}{\delta\,(1-\tfrac{\delta}{2})^{\tau}}$, $F_{0}\geq f(\mathbf{x}_{0})-f^{*}$.
Algorithm Non-convex $\mu$-strongly convex
$D$-EF-SGD $\begin{aligned} \mathcal{O}(&\frac{\sigma^{2}}{n\,\epsilon^{2}}+\frac{\sqrt{\;\tfrac{\phi\,\zeta^{2}}{\delta}+\left(\phi+\tfrac{\tau}{n}\right)\sigma^{2}\;}}{\epsilon^{3/2}}+\\[-3.0pt] &\frac{\tau+Z\sqrt{\phi/\delta}}{\epsilon})\cdot LF_{0}\end{aligned}$ $\begin{aligned} \mathcal{O}(&\frac{\sigma^{2}}{n\,\mu\,\epsilon}+\frac{\sqrt{\,L\!\left(\tfrac{\phi\,\zeta^{2}}{\delta}+\left(\phi+\tfrac{\tau}{n}\right)\sigma^{2}\right)\,}}{\mu\,\epsilon^{1/2}}+\\[-3.0pt] &\frac{L(\tau+Z\sqrt{\phi/\delta})}{\mu})\end{aligned}$
D-SGD with delayed aggregation DGA $\mathcal{O}(\frac{\sigma^{2}}{n\,\epsilon^{2}}+\frac{\sqrt{\;\tfrac{\tau}{n}\,\sigma^{2}\;}}{\epsilon^{3/2}}+\frac{\tau}{\epsilon})\cdot LF_{0}$ $\mathcal{O}(\frac{\sigma^{2}}{n\,\mu\,\epsilon}+\frac{\sqrt{\,L\,\tfrac{\tau}{n}\,\sigma^{2}\,}}{\mu\,\epsilon^{1/2}}+\frac{L\tau}{\mu})$
D-SGD with gradient compression stich2020communication $\mathcal{O}(\frac{\sigma^{2}}{n\,\epsilon^{2}}+\frac{\sqrt{\;\tfrac{\zeta^{2}}{\delta}+\sigma^{2}\;}}{\sqrt{\delta}\,\epsilon^{3/2}}+\frac{Z}{\delta\,\epsilon})\cdot LF_{0}$ $\mathcal{O}(\frac{\sigma^{2}}{n\,\mu\,\epsilon}+\frac{\sqrt{\,L\!\left(\tfrac{\zeta^{2}}{\delta}+\sigma^{2}\right)\,}}{\sqrt{\delta}\,\mu\,\epsilon^{1/2}}+\frac{LZ}{\mu\,\delta})$
@@
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth. There exists a stepsize $\gamma\leq\min\{\frac{1}{4L\tau},\frac{1}{4LZ\sqrt{\phi/\delta}}\}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, such that at most
$\begin{split}\mathcal{O}\Bigg(&\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\tfrac{\phi\zeta^{2}}{\delta}+\left(\phi+\tfrac{\tau}{n}\right)\sigma^{2}}}{\epsilon^{3/2}}+\frac{\tau}{\epsilon}+\frac{Z\sqrt{\phi}}{\sqrt{\delta}\,\epsilon}\Bigg)\cdot L\,(f(\mathbf{x}_{0})-f^{*})\end{split}$
@@
Remark 1 ($\phi$ determining the convergence in the non-degradation condition).
In typical distributed training scenarios, the number of workers $n$ is large and $\frac{\tau}{n}\leq 1$. As the training scale increases, $\tau$ tends to decrease due to the larger model requiring larger computation time, and $n$ tends to increase, resulting in an even smaller $\frac{\tau}{n}$. While to reduce the communication traffic, $\delta$ tends to be $10\%$ or less, common in training with gradient compression dgc; grace, leading to the order of magnitude of $\phi$ larger than that of $\frac{\tau}{n}$. Then the convergence rate can be written as $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}+\frac{Z\sqrt{\frac{\phi}{\delta}}+\tau}{\epsilon})$. Similarly to previous work dagctmc, the frist term $\frac{\sigma^{2}}{n\epsilon^{2}}$ is independent of $\delta$ and $\tau$. The second term $\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}$ has a greater impact on the overall convergence rate compared to the third term.
Let us focus on $\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}$, in LLMs training scenarios (like pre-training or full-parameter fine-tuning), local datasets are typically allocated centrally, resulting in IID datasets ($\zeta\approx 0$ ). Moreover, due to the small batch sizes commonly used in LLM training (limited by the constraint of GPU memory), $\sigma$ tends to be large. Under such circumstances, $\frac{\zeta^{2}}{\delta}\ll\sigma^{2}$ and $\phi$ captures the convergence rate. While in non-IID scenarios, (large $\zeta$) or training small models (small $\sigma^{2}$), the key factor is $\phi^{\prime}=\frac{1-\delta}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}$. In this work, we focus on the first scenario where $\phi$ is the key factor that determines the convergence of DD-EF-SGD. For the analysis in non-IID scenarios (like Federated Learning), we use $\phi^{\prime}$ as the key factor.
Remark 2 (Degradation condition).
DD-EF-SGD has two degradation conditions, namely, without compression ($\delta=1$) and without delayed aggregation ($\tau=0$). When $\delta=1$, $\phi=0$ and DD-EF-SGD degrades into D-SGD with delayed aggregation. Then the convergence rate is equal to $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\frac{\tau}{n}\sigma^{2}}}{\epsilon^{3/2}}+\frac{\tau}{\epsilon})$, same with the convergence rate of D-SGD with delayed aggregation stich2022sharper; stich2019error. When $\tau=0$, DD-EF-SGD degrades to D-SGD with gradient compression, and $\phi=\frac{1-\delta}{\delta}$. Then the convergence rate becomes $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\zeta/\sqrt{\delta}+\sigma}{\sqrt{\delta}\epsilon^{3/2}}+\frac{Z}{\delta\epsilon})$, same with the convergence rate of D-SGD with gradient compression stich2020communication.
Theorem 2 (Convex convergence rate of DD-EF-SGD).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-convex. Then there exists a stepsize $\gamma\leq\min\{\frac{1}{4\sqrt{2}L\tau},\frac{1}{16LZ\sqrt{2\phi/\delta}}\}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L\tau}{\mu}+\frac{LZ\sqrt{\phi/\delta}}{\mu})\end{split}$
iterations of DD-EF-SGD, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{\mathbf{T}-1}\right\}$, selected probabilistically based on $(1-\frac{\mu\gamma}{2})^{-t}$.
@@
Same to remarks in Theorem 1, in the degradation conditions, the convex convergence rate is the same as that of D-SGD with delayed aggregation (under $\delta=1$) and D-SGD with gradient compression (under $\tau=0$). In the non-degradation conditions, the convergence rate can be written into $\mathcal{O}(\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L(\tau+Z\sqrt{\phi/\delta})}{\mu})$, $\phi$ dominates the convergence rate of DD-EF-SGD in the convex case, too.